### v1

# On the Convergence of HalpernSGD

Abstract
HalpernSGD is a gradient-type optimizer obtained by combining the classical Halpern fixed-point iteration with a stochastic gradient step. While its empirical advantages have already been observed in [12, 16], this paper provides a theoretical analysis of the method. Assuming a convex $L$-smooth objective and standard stochastic-approximation conditions , we prove almost sure convergence of the iterates to a minimizer, characterized as the metric projection of the anchor onto the solution set. These results fill a gap in the literature on HalpernSGD and lay the groundwork for future extensions to adaptive schemes.
keywords:
Halpern iteration; stochastic gradient methods; nonexpansive operators; convex optimization; almost sure convergence.
1 Introduction
One of the main difficulties in training deep neural networks is the design of effective strategies to reduce the loss. From a mathematical viewpoint, this is typically modeled as an unconstrained optimization problem of the form
$\min_{x\in\mathbb{R}^{d}}f(x),$
where $x$ denotes the collection of network parameters (weights and biases) and $f$ is the training objective (loss) defined on a $d$-dimensional parameter space.
Since modern architectures are typically highly over-parameterized, an optimizer should be easy to implement, memory-light, and able to scale to large dimensions without a substantial loss of efficiency. For these reasons, first-order methods remain the standard choice in practice, as gradients can be computed efficiently via backpropagation [31]. In particular, in the literature, widely used optimizers can be broadly grouped into two families. The first is the gradient-based family, which uses a global learning rate shared by all coordinates and includes variants such as standard SGD [30], Nesterov accelerated gradient (NAG) [25], Stochastic Gradient Descent (SGD) with momentum [32], and the heavy-ball method (HB) [27]. These methods are often regarded as strong baselines and are widely used in vision pipelines such as image classification [20], object detection [29], and semantic segmentation [1].
The second is the adaptive family, where learning rates are chosen coordinate-wise (or element-wise), including Adam [22], RMSProp [18], AdaGrad [14], and AdaDelta [37]. Adaptive schemes are commonly preferred in settings where rapid progress during training is crucial, e.g., GANs [17], transformers [33], and deep reinforcement learning [24]. At the same time, it is by now well documented that the convergence behavior of Adam-type methods can be subtle: counterexamples and non-convergence regimes are discussed in [28, 38], and oscillatory phenomena (including $2$-limit cycles) are reported in [6]. Moreover, in [16, 12], we comment on an imprecision in the original Adam convergence analysis, specifically in Lemma 10.3 of [22].
A different viewpoint is provided by Metric Fixed Point Theory, which studies iterative schemes for approximating fixed points of nonexpansive and averaged operators; see, e.g., [11, 4, 5]. This framework is particularly appealing because it naturally accommodates infinite-dimensional settings and yields convergence results that do not depend on the dimension of the underlying space. The connection with gradient methods is classically explained through the Baillon–Haddad theorem [2, 3]: for convex $L$-smooth objectives, the forward (gradient) operator is averaged nonexpansive, and its fixed points coincide with minimizers of the loss. Consequently, Gradient Descent (GD),
$x_{n+1}=x_{n}-\eta\nabla f(x_{n}),$ (1)
can be interpreted as a Krasnosel’skiĭ–Mann type fixed-point iteration, for which the convergence speed is often expressed via asymptotic regularity (operator residuals). In particular, the classical KM dynamics yields the residual decay $\|x_{n}-T(x_{n})\|=\mathcal{O}(n^{-1/2})$ as $n\to\infty$ (where $n$ denotes the iteration index), and improved residual rates can be obtained, for instance, through accelerated Krasnosel’skiĭ–Mann dynamics [7].
Within the same fixed-point paradigm, the Halpern iteration is known to exhibit a faster asymptotic-regularity behavior under standard control conditions achieving the sharper rate $\mathcal{O}(n^{-1})$, and it has attracted notable interest in optimization; see, e.g., [13, 23, 36]. Motivated by this phenomenon, we consider an anchored scheme obtained by blending the Halpern iteration with the Gradient Descent (GD) step. In its deterministic form (HalpernGD), the update reads
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(x_{n}-\eta\nabla f(x_{n})\bigr),$ (2)
where $u$ is fixed and termed anchor and $\{\alpha_{n}\}\subset(0,1)$ is a control sequence. When $f$ is convex and $L$-smooth and $\eta\in(0,2/L)$, the associated gradient operator is averaged nonexpansive, so Halpern-type convergence guarantees extend naturally to (2).
From a practical standpoint, however, full-gradient steps such as (1) are often prohibitive in large-scale learning. Mini-batch gradients provide nearly the same descent direction at a fraction of the cost, and small batches may even act as an implicit regularizer [21]. This leads to the stochastic Halpern-type iteration (HalpernSGD), which can be written as
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(x_{n}-\epsilon_{n}(\nabla f(x_{n})+U_{n})\bigr),$ (3)
where $\{U_{n}\}$ is a martingale difference sequence known as noise and $\epsilon_{n}\downarrow 0$ is a diminishing step size. This optimizer, HalpernSGD, was proposed and experimentally investigated in our earlier work [16, 12], where we observed that the anchoring mechanism can substantially reduce the number of training epochs needed to reach a given accuracy. Additional qualitative experiments on standard nonconvex benchmarks (Rastrigin and Rosenbrock functions) further illustrate distinctive optimization trajectories and stability patterns when compared to SGD and Adam.
The purpose of this paper is to provide a rigorous convergence analysis of HalpernSGD (3) under a classical convex $L$-smooth setting and standard stochastic assumptions. Here we establish almost sure convergence of the iterates toward a distinguished minimizer characterized as the projection of the anchor $u$ onto the solution set. Our analysis relies on a suitable coupling between $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$ and on control conditions in the spirit of [35], which are natural in Halpern-type schemes. These results fill a gap in the current literature on HalpernSGD optimizer and provide a foundation for future developments, including the systematic study of anchored variants of adaptive optimizers.
Section 3 collects notation and assumptions, Section 4 contains the boundedness lemma and the almost sure convergence theorem and Section 5 concludes the paper and briefly outlines directions for future work.
2 Related Work and Contributions
The Halpern iteration originated as a fixed-point method for nonexpansive mappings and, under standard control conditions on the weight sequence, guarantees strong convergence toward the fixed point selected by the anchor, typically the fixed point closest to the anchor, see e.g. [19, 34, 35].
In the context of monotone inclusions and variational inequalities, [13] shows how variants of the Halpern iteration can produce near-optimal and parameter-free methods for standard classes of problems with Lipschitz monotone operators.
These initial results are formulated in an essentially deterministic framework; as for the stochastic case, the available works tend to impose variance control, obtained often via increasing mini-batches or variance reduction, to prevent error accumulation.
A starting point is [10], which studies stochastic monotone inclusions and combines variants of Halpern with a recursive PAGE-type estimator to obtain variance that decreases as $O(\varepsilon^{2}/k)$, through data-driven choices of the batch-sizes. The work assumes:
  1. 1.
    unbiasedness,
  2. 2.
    bounded variance,
  3. 3.
    a multi-point oracle assumption that allows querying the oracle at multiple points using the same seed, and
  4. 4.
    a Lipschitz property in expectation.
This line therefore puts at its core the idea that, with constant-variance noise, additional mechanisms are needed to make the effective error decay [10].
A related perspective, but more related to fixed-point theory, is developed in [8], which analyzes a stochastic Halpern iteration in normed spaces and discusses applications in reinforcement learning. Here, the stochastic oracle is modeled as a random approximation $\widetilde{T}(x,\xi)$ of a nonexpansive operator $T$ and the almost sure convergence to a fixed point under the undesiderable hypothesis that $\sum_{n}\|U_{n}\|<+\infty,$ where $\{U_{n}\}$ is the error sequence (Theorem 3.10).
More recently, [26] proposes a general framework of stochastic Halpern scheme for nonexpansive operators, where stochasticity enters as additive perturbations $\xi_{n},\delta_{n}.$ Here, the primary objective is to establish the asymptotic regularity of the sequence of iterates, that is, the convergence $|x_{n+1}-x_{n}|\to 0$. This type of result is obtained under standard conditions on $\alpha_{n}$ and $\beta_{n}$, together with the convergence of $\sum_{n}\mathbb{E}|\delta_{n+1}-\delta_{n}|$ and of $\sum_{n}\mathbb{E}|\xi_{n+1}-\xi_{n}|$.
Alongside the works specifically focused on Halpern-type schemes, general results on stochastic fixed-point iterations are also relevant. In particular, [8] analyzes Stochastic Krasnosel’skiĭ–Mann (SKM) iterations for nonexpansive operators, where the noise sequence is modeled as additive stochastic perturbations with zero mean and finite second moment, that is, $\mathbb{E}[U_{n}\mid\mathcal{F}_{n-1}]=0$ and $\mathbb{E}|U_{n}|^{2}<\infty$. Under standard stepsize assumptions, the authors establish almost sure convergence of the SKM iterates to a fixed point of $T$, without requiring the variance to be uniformly bounded.
In our work we place ourselves in this latter perspective. Indeed, we study HalpernSGD as an “anchored” scheme for $L$-smooth convex minimization, modeling gradients stochastically perturbed by a unbiased noise. Almost sure convergence in our setting is established in a different way than in the works reviewed above. Specifically, we study almost sure convergence to a minimization point, that is, to a zero of the gradient of the objective function $f$, by introducing a second control sequence $\{\varepsilon_{n}\}$ that plays the role of a learning rate. This additional degree of control allows us to impose only a mild condition on the noise, namely the summability condition
$\sum_{n}\varepsilon_{n}^{2}\,\|U_{n}\|^{2}<+\infty,$
which ensures that the cumulative effect of the stochastic perturbations remains tractable. As a consequence, we are able to handle almost sure convergence of the entire sequence $\{x_{n}\}$ even in the presence of noise with merely bounded variance. This, in turn, allows the Halpern framework to be applied in settings where variance reduction or vanishing-noise assumptions are not available.
By exploiting a quadratic-mean recurrence and lemma proved by Xu in [35], we first prove convergence in $L^{2}$ and then almost sure convergence of the whole sequence toward the minimizer selected by the anchor, $x^{\star}=P_{S}(u)$.
In summary, compared to works that require variance reduction to obtain convergence in the stochastic Halpern case [10, 8, 26], our contribution aims at an almost sure guarantee of convergence under bounded-variance assumptions and is based on an appropriate choice and coupling of two control sequences that make the cumulative effect of the noise tractable.
3 Preliminaries
Throughout, $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ denote the Euclidean inner product and norm on $\mathbb{R}^{d}$. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and differentiable with $L$-Lipschitz continuous gradient, i.e.,
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\qquad\forall x,y\in\mathbb{R}^{d}.$ (4)
We denote the set of minimizers by
$S:=\arg\min_{x\in\mathbb{R}^{d}}f(x),$ (5)
and assume $S\neq\emptyset$. Given an anchor $u\in\mathbb{R}^{d}$, we write $P_{S}(u)$ for the (unique) metric projection of $u$ onto $S$ (recall that $S$ is closed and convex for convex $f$).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $\{\mathcal{F}_{n}\}_{n\geq 0}$ be a filtration such that $x_{n}$ (see Equation (3)) is $\mathcal{F}_{n}$-measurable. We model stochastic gradients through a martingale-difference noise sequence $\{U_{n}\}_{n\geq 0}$ satisfying
$\mathbb{E}[\,U_{n}\mid\mathcal{F}_{n}\,]=0\qquad\text{a.s.}\quad\forall n\geq 0,$ (6)
together with a bounded conditional second moment:
$\mathbb{E}\bigl[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\bigr]\leq\sigma^{2}\qquad\text{a.s.}\quad\forall n\geq 0,$ (7)
for some $\sigma\geq 0$. For stepsizes $\{\epsilon_{n}\}_{n\geq 0}\subset(0,\infty)$ and Halpern weights $\{\alpha_{n}\}_{n\geq 0}\subset(0,1)$, the stochastic forward (gradient) operator is
$\widetilde{T}_{n}(x):=x-\epsilon_{n}\bigl(\nabla f(x)+U_{n}\bigr),$ (8)
and the HalpernSGD iteration reads
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\,\widetilde{T}_{n}(x_{n})=\alpha_{n}u+(1-\alpha_{n})\Bigl(x_{n}-\epsilon_{n}\bigl(\nabla f(x_{n})+U_{n}\bigr)\Bigr).$ (9)
When $U_{n}\equiv 0$, (9) reduces to the deterministic Halpern iterative scheme.
In this context, a decreasing stepsize sequence $\{\epsilon_{n}\}$ is required to balance progress and noise reduction when moving from deterministic to stochastic gradients. Classical stochastic approximation theory [30] prescribes the following conditions:
$\sum_{n=1}^{\infty}\epsilon_{n}=\infty,\qquad\sum_{n=1}^{\infty}\epsilon_{n}^{2}<\infty.$ (10)
As it will be specified in the next section, in our analysis, $\{\epsilon_{n}\}$ will be additionally coupled with $\{\alpha_{n}\}$ through Halpern-type control conditions.
We recall a few standard facts that will be used in the next section. Convexity of $f$ implies monotonicity of the gradient:
$\langle\nabla f(x)-\nabla f(y),\,x-y\rangle\geq 0\qquad\forall x,y\in\mathbb{R}^{d}.$ (11)
Moreover, the $L$-smoothness assumption (4) together with convexity yields the Baillon–Haddad cocoercivity inequality
$\langle\nabla f(x)-\nabla f(y),\,x-y\rangle\geq\frac{1}{L}\,\|\nabla f(x)-\nabla f(y)\|^{2}\qquad\forall x,y\in\mathbb{R}^{d}.$ (12)
In particular, if $x^{*}\in S$, then $\nabla f(x^{*})=0$ and (12) gives
$\langle\nabla f(x),\,x-x^{*}\rangle\geq\frac{1}{L}\,\|\nabla f(x)\|^{2}\qquad\forall x\in\mathbb{R}^{d}.$ (13)
We also recall the standard gradient-norm bound (sometimes called a “Polyak–Łojasiewicz-type inequality” for smooth convex functions):
$\frac{1}{2L}\,\|\nabla f(x)\|^{2}\leq f(x)-f^{*}\qquad\forall x\in\mathbb{R}^{d},$ (14)
where $f^{*}:=\min_{\mathbb{R}^{d}}f$. Finally, since $f$ is convex and differentiable, minimizers are characterized by the first-order condition
$x^{*}\in S\quad\Longleftrightarrow\quad\nabla f(x^{*})=0.$ (15)
4 Convergence Analysis of HalpernSGD
We now investigate the convergence properties of the HalpernSGD iteration (9). Throughout this section, we consider the standard stochastic setting where $\{\mathcal{F}_{n}\}$ denotes the natural filtration generated by the iterates and the random perturbations, i.e.
$\mathcal{F}_{n}:=\sigma(x_{0},U_{0},\dots,U_{n-1}).$
Following the standard setting used in stochastic fixed-point theory (see, e.g.,[9]), the perturbations $U_{n}$ are assumed to satisfy the martingale difference conditions:
$\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0,\qquad\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\leq\sigma^{2}\quad\text{a.s. for all }n\in\mathbb{N},$
for some constant $\sigma>0$. The stepsizes $\{\epsilon_{n}\}$ and Halpern weights $\{\alpha_{n}\}$ are real sequences satisfying
$\epsilon_{n}>0,\quad\epsilon_{n}\to 0,\quad\sum_{n=0}^{\infty}\epsilon_{n}=+\infty,\quad\sum_{n=0}^{\infty}\epsilon_{n}^{2}<+\infty,$
and
$\alpha_{n}\in(0,1),\quad\alpha_{n}\to 0,\quad\sum_{n=0}^{\infty}\alpha_{n}=+\infty,\quad\sum_{n=0}^{\infty}\alpha_{n}^{2}<+\infty$
with
$\lim_{n\rightarrow+\infty}\frac{\epsilon_{n}}{\alpha_{n}}=+\infty\quad\text{and }\sum_{n=0}^{\infty}\alpha_{n}\epsilon_{n}<\infty.$
Lemma 4.1.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a convex and $L$-smooth function, and let $\{x_{n}\}$ be the sequence generated by the HalpernSGD iteration (9). Let $x^{*}\in S$ (equivalently, $\nabla f(x^{*})=0$) and assume that the stochastic perturbations and the parameters $\{\alpha_{n}\},\{\epsilon_{n}\}$ satisfy the conditions stated above, and in addition that $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$. Then, the sequence $\{\mathbb{E}\|x_{n}-x^{*}\|^{2}\}_{n\in\mathbb{N}}$ is bounded.
Proof.
Let $x^{*}$ be a fixed point of $T$, and consequently $\nabla f(x^{*})=0$. By definition of the HalpernSGD iteration, we have
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\big(x_{n}-\epsilon_{n}(\nabla f(x_{n})+U_{n})\big),$
which yields
$\|x_{n+1}-x^{*}\|^{2}=\big\|\alpha_{n}(u-x^{*})+(1-\alpha_{n})\big(x_{n}-x^{*}-\epsilon_{n}(\nabla f(x_{n})+U_{n})\big)\big\|^{2}.$
Let us denote $g_{n}(x_{n})=\nabla f(x_{n})+U_{n}$. Using the convexity of the squared norm, we obtain
$\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\|x_{n}-x^{*}-\epsilon_{n}g_{n}(x_{n})\|^{2},$
and expanding the second term gives
$\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\!\left[\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\|g_{n}(x_{n})\|^{2}-2\epsilon_{n}\langle g_{n}(x_{n}),x_{n}-x^{*}\rangle\right].$
Taking the conditional expectation with respect to $\mathcal{F}_{n}$ yields
$\mathbb{E}\!\left[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}\right]\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})\epsilon_{n}\langle\mathbb{E}[g_{n}(x_{n})\mid\mathcal{F}_{n}],x_{n}-x^{*}\rangle$
$(1-\alpha_{n})\epsilon_{n}^{2}\mathbb{E}\!\left[\|g_{n}(x_{n})\|^{2}\mid\mathcal{F}_{n}\right].$
We analyze each term separately.
  1. 1.
    Since $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$, we have $\mathbb{E}[g_{n}(x_{n})\mid\mathcal{F}_{n}]=\nabla f(x_{n})$. Using convexity of $f$, it follows that
    $\langle\nabla f(x_{n})-\nabla f(x^{*}),x_{n}-x^{*}\rangle\geq 0,$
    so that the cross term is nonpositive:
    $2(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle\leq 0.$
  2. 2.
    By the properties of conditional expectation,
    $\mathbb{E}\big[\mathbb{E}[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}]\big]=\mathbb{E}\|x_{n+1}-x^{*}\|^{2}.$
  3. 3.
    Using $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$ again, we have that
    $\mathbb{E}\left[\langle U_{n},\nabla f(x_{n})\rangle\right]=\mathbb{E}\Big[\mathbb{E}\big[\langle U_{n},\nabla f(x_{n})\rangle\mid\mathcal{F}_{n}\big]\Big]=\mathbb{E}\Big[\langle\mathbb{E}\big[U_{n}\mid\mathcal{F}_{n}\big],\nabla f(x_{n})\rangle\Big]=0\text{ and}$
    $\mathbb{E}\|g_{n}(x_{n})\|^{2}=\mathbb{E}\|\nabla f(x_{n})+U_{n}-\nabla f(x^{*})\|^{2}=\mathbb{E}\|\nabla f(x_{n})-\nabla f(x^{*})\|^{2}+\mathbb{E}\|U_{n}\|^{2}.$
    Since $f$ is $L$-smooth, $\|\nabla f(x_{n})-\nabla f(x^{*})\|\leq L\|x_{n}-x^{*}\|$, hence
    $\mathbb{E}\|g_{n}(x_{n})\|^{2}\leq L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\mathbb{E}\|U_{n}\|^{2}.$
Combining these estimates and taking expectations, we obtain
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\bigl(\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\mathbb{E}\|U_{n}\|^{2}\bigr)$
$\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\mathbb{E}\|U_{n}\|^{2}.$
Let $a_{n}:=\mathbb{E}\|x_{n}-x^{*}\|^{2}$ and $\mu_{n}:=\mathbb{E}\|U_{n}\|^{2}$. Then
$a_{n+1}\leq\alpha_{n}\|u-x^{*}\|^{2}+\big((1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\big)a_{n}+\epsilon_{n}^{2}\mu_{n}.$
$(\alpha_{n}-\epsilon_{n}^{2}L^{2})\|u-x^{*}\|^{2}+\big[1-(\alpha_{n}-\epsilon_{n}^{2}L^{2})\big]a_{n}+\epsilon_{n}^{2}\mu_{n}+\epsilon_{n}^{2}L^{2}\|u-x^{*}\|^{2}.$
Setting $\gamma_{n}:=\alpha_{n}-\epsilon_{n}^{2}L^{2}$, we can rewrite
$a_{n+1}\leq\gamma_{n}\|u-x^{*}\|^{2}+(1-\gamma_{n})a_{n}+\epsilon_{n}^{2}(\mu_{n}+L^{2}\|u-x^{*}\|^{2}).$
We remark that $\gamma_{n}\in(0,1)$. Indeed, $\gamma_{n}=\alpha_{n}-\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, and by the coupling condition $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}$, we have $\gamma_{n}>0$. Then, iterating the inequality gives
$a_{n+1}\leq\max\{\|u-x^{*}\|^{2},a_{0}\}+\sum_{k=0}^{n}\epsilon_{k}^{2}\big[\mu_{k}+L^{2}\|u-x^{*}\|^{2}\big].$
Now, by assumptions, $\sum_{k}\epsilon_{k}^{2}<\infty$ and the stochastic perturbations satisfy $\mathbb{E}[\|U_{k}\|^{2}\mid\mathcal{F}_{k}]\leq\sigma^{2}$ a.s. $\forall k\in\mathbb{N}$, then $\mu_{k}\leq\sigma^{2}$ and
$\sum_{k=0}^{\infty}\epsilon_{k}^{2}(\mu_{k}+L^{2}\|u-x^{*}\|^{2})\leq(\sigma^{2}+L^{2}\|u-x^{*}\|^{2})\sum_{k=0}^{\infty}\epsilon_{k}^{2}<\infty.$
Hence, the sequence $\{a_{n}\}$ is bounded, that is, $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty.\qed$
Remark 1.
From a practical standpoint, in view of the experimental analysis, the condition $\alpha_{n}-\epsilon_{n}^{2}L^{2}>0$ can be simplified through a sufficient asymptotic requirement. Indeed, by assuming
$\lim_{n\to+\infty}\frac{\alpha_{n}}{\epsilon_{n}^{2}}=+\infty,$ (16)
it follows that there exists $\tilde{n}\in\mathbb{N}$ such that
$\forall n>\tilde{n},\qquad\frac{\alpha_{n}}{\epsilon_{n}^{2}}>L^{2},$
which directly ensures $\alpha_{n}-\epsilon_{n}^{2}L^{2}>0$ for all sufficiently large $n$. This condition is easier to verify in numerical implementations, where both sequences $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$ are explicitly specified.
Theorem 4.2.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth, $x^{*}:=P_{S}(u)$ and $\{x_{n}\}$ be generated by (9). Assume the stochastic setting and the step-size conditions stated above, together with $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$.
Then
$x_{n}\xrightarrow{\text{\;a.s.}}x^{*},$
i.e., $x^{*}$ is the unique point in $S$ satisfying the variational inequality
$\langle u-x^{*},\,p-x^{*}\rangle\leq 0\qquad\text{for all }p\in S.$
Proof.
We show that $\mathbb{E}\|\nabla f(x_{n})\|^{2}\to 0$ under the stated assumptions. By definition of HalpernSGD iteration, we have
$\|x_{n+1}-x^{*}\|^{2}=\bigl\|\alpha_{n}(u-x^{*})+(1-\alpha_{n})\bigl(x_{n}-x^{*}-\epsilon_{n}g_{n}\bigr)\bigr\|^{2}\qquad g_{n}:=\nabla f(x_{n})+U_{n}.$
Expanding the square we obtain:
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,x_{n}-x^{*}-\epsilon_{n}g_{n}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}-\epsilon_{n}g_{n}\|^{2}=$
$\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle g_{n},u-x^{*}\rangle$
$(1-\alpha_{n})^{2}\bigl(\|x_{n}-x^{*}\|^{2}-2\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+\epsilon_{n}^{2}\|g_{n}\|^{2}\bigr).$
Collecting terms, we can rewrite this as
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle g_{n},u-x^{*}\rangle$
$(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})^{2}\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\|g_{n}\|^{2}.$
Now, taking the conditional expectation with respect to $\mathcal{F}_{n}$ and using $\mathbb{E}[g_{n}\mid\mathcal{F}_{n}]=\nabla f(x_{n})$ (since $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$), we obtain
$\mathbb{E}\!\left[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}\right]$ $=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle$
$\quad-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),u-x^{*}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}$
$\quad-2(1-\alpha_{n})^{2}\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\,\mathbb{E}\!\left[\|g_{n}\|^{2}\mid\mathcal{F}_{n}\right].$
First, since $f$ is convex and $L$-smooth, by the co-coercivity inequality we have
$2(1-\alpha_{n})^{2}\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle\;\leq\;-\frac{2(1-\alpha_{n})^{2}\epsilon_{n}}{L}\,\|\nabla f(x_{n})\|^{2}.$
Next, from Lemma 4.1, the sequence of iterates is bounded in mean square, that is, there exists a constant $C>0$ such that $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\leq C$. Thus, by applying the Cauchy–Schwarz inequality and the $L$-Lipschitz continuity of $\nabla f$, we have $|\langle\nabla f(x_{n}),u-x^{*}\rangle|\leq\|\nabla f(x_{n})\|\,\|u-x^{*}\|\leq L\sqrt{C}\,\|u-x^{*}\|.$ Hence,
$2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),u-x^{*}\rangle\leq 2\alpha_{n}(1-\alpha_{n})\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
Finally, by the martingale-difference assumption, $\mathbb{E}\langle U_{n},\nabla f(x_{n})\rangle=\mathbb{E}\big[\langle\mathbb{E}[U_{n}\mid\mathcal{F}_{n}],\,\nabla f(x_{n})\rangle\big]=0,$ and then, by bounded conditional variance assumption, we have
$\mathbb{E}\!\left[\|g_{n}\|^{2}\mid\mathcal{F}_{n}\right]=\|\nabla f(x_{n})\|^{2}+\mathbb{E}\!\left[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\right]\leq\|\nabla f(x_{n})\|^{2}+\sigma^{2}.$
Then, taking the total expectation and using $(1-\alpha_{n})\leq 1$, we obtain
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$ $\leq\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}-(1-\alpha_{n})^{2}\!\Bigl(\tfrac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\,\mathbb{E}\|\nabla f(x_{n})\|^{2}$
$\quad+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$ (17)
By the coupling $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ we have $\epsilon_{n}\leq 1/L$ for all sufficiently large $n$, whence $\Bigl(\tfrac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\ \geq\ \tfrac{\epsilon_{n}}{L}.$ Thus, setting $\beta_{n}:=\frac{(1-\alpha_{n})^{2}\epsilon_{n}}{L\alpha_{n}}>0$, (17) implies
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}-\mathbb{E}\|x_{n}-x^{*}\|^{2}$ $\leq\underbrace{\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|}_{=:R_{n}}$
$\quad+\alpha_{n}\,\underbrace{\Bigl(2\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle-\beta_{n}\,\mathbb{E}\|\nabla f(x_{n})\|^{2}\Bigr)}_{=:\,\delta_{n}}.$
With the definitions of $R_{n}$ and $\delta_{n}$, the previous inequality becomes
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}-\mathbb{E}\|x_{n}-x^{*}\|^{2}\;\leq\;R_{n}+\alpha_{n}\,\delta_{n}.$ (18)
By the step-size assumptions, $\sum_{n}\alpha_{n}^{2}<\infty$, $\sum_{n}\epsilon_{n}^{2}<\infty$ and $\sum_{n}\alpha_{n}\epsilon_{n}<\infty$, hence $\sum_{n}R_{n}<\infty$. Summing (18) from $n=0$ to N1N-1 and using the telescoping of the left-hand side, we get
$\sum_{n=0}^{N-1}\alpha_{n}\delta_{n}\ \geq\ \mathbb{E}\|x_{N}-x^{*}\|^{2}-\mathbb{E}\|x_{0}-x^{*}\|^{2}-\sum_{n=0}^{N-1}R_{n}\ \geq\ -\,\Bigl(\mathbb{E}\|x_{0}-x^{*}\|^{2}+\sum_{n=0}^{\infty}R_{n}\Bigr)=:-\,\widetilde{M}$,
for all $N$, with some finite constant $\widetilde{M}$$\infty$ independent of $N$. Since $\sum_{n=0}^{N-1}\alpha_{n}\delta_{n}$ is bounded from below and $\sum_{n}\alpha_{n}=+\infty$, it follows that a null subsequence $\{\delta_{n_{k}}\}$ must exist, i.e.
$\lim_{k\to 0}\Bigl(2\langle u-x^{*},\,\mathbb{E}[x_{n_{k}}-x^{*}]\rangle-\beta_{n_{k}}\,\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\Bigr)=0.$ (19)
Observe that Lemma 4.1 implies that $2\langle u-x^{*},\,\mathbb{E}[x_{n_{k}}-x^{*}]\rangle$ remains bounded, while $\beta_{n_{k}}\sim\epsilon_{n_{k}}/\alpha_{n_{k}}$ diverges.
These facts readily implies that $\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\to 0$ and, by eventually passing to further subsequences, that
$\nabla f(x_{n_{k}})\to 0\text{ a.s.}$ (20)
as well.
Now, observe that for every $n$,
$\mathbb{E}\bigl[\|x_{n+1}-x_{n}\|^{2}\,\big|\,\mathcal{F}_{n}\bigr]=\mathbb{E}\bigl[\bigl\|\alpha_{n}(u-x_{n})+(1-\alpha_{n})\bigl(-\epsilon_{n}(\nabla f(x_{n})+U_{n}\bigr))\bigr\|^{2}\big|\,\mathcal{F}_{n}\bigr].$
Using the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$, we obtain
$\mathbb{E}\bigl[\|x_{n+1}-x_{n}\|^{2}\,\big|\,\mathcal{F}_{n}\bigr]\;\leq\;2\alpha_{n}^{2}\|u-x_{n}\|^{2}+2\epsilon_{n}^{2}\bigl(\|\nabla f(x_{n})\|^{2}+\sigma^{2}\bigr).$
Since $\mathbb{E}\|u-x_{n}\|^{2}$ and $\mathbb{E}\|\nabla f(x_{n})\|^{2}$ are bounded, by Lemma 4.1, while $\sum_{n}\alpha_{n}^{2}<\infty$ and $\sum_{n}\epsilon_{n}^{2}<\infty$, taking total expectations we deduce that
$\sum_{n=0}^{\infty}\mathbb{E}\|x_{n+1}-x_{n}\|^{2}<\infty,$
Hence, by a standard convergence argument (see, e.g., [15]), it follows that
$\|x_{n+1}-x_{n}\|\;\to\;0\quad\text{almost surely.}$
Moreover, since $\nabla f$ is $L$-Lipschitz, $\|\nabla f(x_{n+1})-\nabla f(x_{n})\|\;\leq\;L\,\|x_{n+1}-x_{n}\|,$ it holds as well
$\|\nabla f(x_{n+1})-\nabla f(x_{n})\|\to 0\text{ a.s.}$
Returning to (20) we can finally conclude that
$\|\nabla f(x_{n})\|\;\to\;0\quad\text{almost surely.}$
Let $(x_{n_{k}})$ be a subsequence such that
[…]
which exists and is finite since the sequence $\bigl(\mathbb{E}\langle x_{n}-x^{*},\,u-x^{*}\rangle\bigr)$ is bounded.
By passing, if necessary, to a further subsequence, we may assume that for almost every $\omega\in\Omega$,
$\langle x_{n_{k}}(\omega)-x^{*},\,u-x^{*}\rangle\;\longrightarrow\;\langle P(\omega)-x^{*},\,u-x^{*}\rangle$
for some cluster point $P(\omega)$ of $(x_{n}(\omega))$.
Since $\nabla f(x_{n_{k}})\to 0$ almost surely, it follows that each limit point $P(\omega)$ satisfies $\nabla f(P(\omega))=0$; that is, $P(\omega)$ is a critical point of $f$. Moreover, since
$\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\|u-x^{*}\|^{2}<\infty,$
the sequence of random variables
$X_{n}:=\langle x_{n}-x^{*},\,u-x^{*}\rangle$
is bounded in $L^{2}$. By the de la Vallée–Poussin criterion, this implies that $\{X_{n}\}$ is uniformly integrable Let $(x_{n_{k}})$ be a subsequence such that
[…]
for some measurable random variable $P$. Then, for those $\omega$,
$X_{n_{k}}(\omega)=\langle x_{n_{k}}(\omega)-x^{*},\,u-x^{*}\rangle\longrightarrow\langle P(\omega)-x^{*},\,u-x^{*}\rangle.$
This last, together with the already proved uniform integrability implies the convergence in $L^{1}$ by means of Vitali’s Theorem. Hence, we deduce that
$\lim_{k\to\infty}\mathbb{E}\,\langle x_{n_{k}}-x^{*},\,u-x^{*}\rangle=\mathbb{E}\,\langle P-x^{*},\,u-x^{*}\rangle.$
Since $\nabla f(P(\omega))=0,$ $P(\omega)\in S$ for almost every $\omega,$ and since $x^{*}=P_{S}(u)$, it follows that
$\langle P(\omega)-x^{*},\,u-x^{*}\rangle\leq 0\quad\text{for almost every }\omega,$
and therefore
$\limsup_{n\to\infty}\mathbb{E}\,\langle x_{n}-x^{*},\,u-x^{*}\rangle\;\leq\;0.$
Returning to (17), we rewrite it as
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$ $\leq(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}+\alpha_{n}^{2}\|u-x^{*}\|^{2}\quad+\epsilon_{n}^{2}\sigma^{2}$
$+2\alpha_{n}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
We observe that the sequences $(\alpha_{n}^{2})$, $(\epsilon_{n}^{2})$, and $(\alpha_{n}\epsilon_{n})$ are summable, while $\limsup_{n\to\infty}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle\leq 0.$ Hence, by applying the Xu Lemma (see [35, Lemma 2.1]), we conclude that
$\mathbb{E}\|x_{n}-x^{*}\|^{2}\;\longrightarrow\;0.$
Consequently, there exists a subsequence $(x_{n_{k}})$ such that $\|x_{n_{k}}-x^{*}\|\to 0$ almost surely. Since we have already established that $\|x_{n+1}-x_{n}\|\to 0$ almost surely, this asymptotic regularity implies that the entire sequence converges:
$x_{n}\;\longrightarrow\;x^{*}\quad\text{almost surely.}$
5 Conclusion and future perspectives
In this work we addressed the convergence theory for the HalpernSGD optimizer. Building on tools from smooth convex analysis and stochastic approximation, we proved that its iterates converge almost surely to a specific solution selected by the anchor, namely $x^{*}=P_{S}(u)$.
As future work, we plan to broaden the experimental analysis carried out in [16, 12], where HalpernSGD was compared mainly with SGD under constant learning rates, in order to assess whether HalpernSGD can be among the most efficient optimizers within the gradient-based family while maintaining (or improving) accuracy. In particular, we aim to benchmark HalpernSGD against SGD [30], SGD with momentum [32], NAG [25], and HB [27] and other accelerated gradient schemes. Motivated by the empirical benefits observed for the anchoring mechanism in [16, 12], we also plan to investigate anchored variants of adaptive methods. A natural direction is to combine an anchor with an Adam-type update through a convex blending (an “Anchored-Adam” scheme), analogously to HalpernSGD as an anchored variant of SGD. Preliminary experiments suggest that Halpern-style anchoring may act as a general acceleration and stabilization mechanism across different optimizer families. Establishing convergence guarantees for such anchored adaptive schemes remains an open theoretical challenge, especially in view of the known subtle convergence behavior of Adam optimizer, and we consider this a key direction for future research.
Funding
This publication was partially funded by the PhD program in Mathematics and Computer Science at University of Calabria, Cycle XXXVIII with the support of a scholarship financed by DM 351/2022 (CUP H23C22000440007), based on the NRPP funded by the European Union.