### v2

# On the Convergence of HalpernSGD

Abstract
HalpernSGD is a gradient-type optimizer obtained by combining the classical Halpern fixed-point iteration with a stochastic gradient step. While its empirical advantages have already been observed in [14, 19], this paper provides a theoretical analysis of the method. Assuming a convex $L$-smooth objective and standard stochastic-approximation conditions, we prove almost sure convergence of the iterates to a minimizer, characterized as the metric projection of the anchor onto the solution set. These results fill a gap in the literature on HalpernSGD and lay the groundwork for future extensions to adaptive schemes.
keywords:
Halpern iteration; stochastic gradient methods; nonexpansive operators; convex optimization; almost sure convergence.
1 Introduction
One of the main difficulties in training deep neural networks is the design of effective strategies to reduce the loss. From a mathematical viewpoint, this is typically modeled as an unconstrained optimization problem of the form
$\min_{x\in\mathbb{R}^{d}}f(x),$
where $x$ denotes the collection of network parameters (weights and biases) and $f$ is the training objective (loss) defined on a $d$-dimensional parameter space.
Since modern architectures are typically highly over-parameterized, an optimizer should be easy to implement, memory-light, and able to scale to large dimensions without a substantial loss of efficiency. For these reasons, first-order methods remain the standard choice in practice, as gradients can be computed efficiently via backpropagation [38]. In particular, in the literature, widely used optimizers can be broadly grouped into two families. The first is the gradient-based family, which uses a global learning rate shared by all coordinates and includes variants such as standard SGD [37], Nesterov accelerated gradient (NAG) [30], Stochastic Gradient Descent (SGD) with momentum [40], and the heavy-ball method (HB) [34]. These methods are often regarded as strong baselines and are widely used in vision pipelines such as image classification [24], object detection [36], and semantic segmentation [2].
The second is the adaptive family, where learning rates are chosen coordinate-wise (or element-wise), including Adam [26], RMSProp [22], AdaGrad [17], and AdaDelta [46]. Adaptive schemes are commonly preferred in settings where rapid progress during training is crucial, e.g., GANs [21], transformers [41], and deep reinforcement learning [29]. At the same time, it is by now well documented that the convergence behavior of Adam-type methods can be subtle: counterexamples and non-convergence regimes are discussed in [35], and oscillatory phenomena (including $2$-limit cycles) are reported in [7]. Moreover, in [19, 14], we comment on an imprecision in the original Adam convergence analysis, specifically in Lemma 10.3 of [26].
A different viewpoint is provided by Metric Fixed Point Theory, which studies iterative schemes for approximating fixed points of nonexpansive and averaged operators; see, e.g., [13, 5, 6]. This framework is particularly appealing because it naturally accommodates infinite-dimensional settings and yields convergence results that do not depend on the dimension of the underlying space. The connection with gradient methods is classically explained through the Baillon–Haddad theorem [3, 4]: for convex $L$-smooth objectives, the forward (gradient) operator is averaged nonexpansive, and its fixed points coincide with minimizers of the loss. Consequently, Gradient Descent (GD),
$x_{n+1}=x_{n}-\eta\nabla f(x_{n}),$ (1)
can be interpreted as a Krasnosel’skiĭ–Mann type fixed-point iteration, for which the convergence speed is often expressed via asymptotic regularity (operator residuals). In particular, the classical KM dynamics yields the residual decay $\|x_{n}-T(x_{n})\|=\mathcal{O}(n^{-1/2})$ as $n\to\infty$ (where $n$ denotes the iteration index), and improved residual rates can be obtained, for instance, through accelerated Krasnosel’skiĭ–Mann dynamics [8].
Within the same fixed-point paradigm, the Halpern iteration is known to exhibit a faster asymptotic-regularity behavior under standard control conditions achieving the sharper rate $\mathcal{O}(n^{-1})$, and it has attracted a particular interest in optimization; see, e.g., [16, 28, 45]. Motivated by this phenomenon, we consider an anchored scheme obtained by blending the Halpern iteration with the Gradient Descent (GD) step. In its deterministic form (HalpernGD), the update reads
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(x_{n}-\eta\nabla f(x_{n})\bigr),$ (2)
where $u$ is fixed and termed anchor and $\{\alpha_{n}\}\subset(0,1)$ is a control sequence. When $f$ is convex and $L$-smooth and $\eta\in(0,2/L)$, the associated gradient operator is averaged nonexpansive, so Halpern-type convergence guarantees extend naturally to (2).
In [20], the authors introduced a variant of the Halpern iteration, where the step size $\eta$ is replaced by a vanishing sequence $\{\epsilon_{n}\}$, and they established strong convergence to a solution under suitable control conditions on $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$. In particular, their results imply the strong convergence of the sequence
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(x_{n}-\epsilon_{n}\partial f(x_{n})\bigr),$ (3)
where $\partial f$ is the subdifferential of a proper, lower semicontinuous, convex function $f$ which can also be unbounded.
Here, the possibly unboundedness of the subdifferential is mitigated by the vanishing step size $\epsilon_{n}$, which prevents the iterates from diverging and allows to establish strong convergence to a minimizer.
From a practical standpoint, however, full-gradient steps such as (2) or (3) are often prohibitive in large-scale learning. Mini-batch gradients provide nearly the same descent direction at a fraction of the cost, and small batches may even act as an implicit regularizer [25]. This leads to the stochastic Halpern-type iteration (HalpernSGD), which can be written as
$X_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl(X_{n}-\epsilon_{n}(\nabla f(X_{n})+U_{n})\bigr),$ (4)
where $\{X_{n}\}$ beacomes a sotchastic process over a probability space $\Omega$, $\{U_{n}\}$ is a martingale difference sequence and $\epsilon_{n}\downarrow 0$ is a vanishing step size.
The remainder of this paper is organized as follows. Section 3 introduces the notation, recalls fundamental properties of smooth convex functions, and states the assumptions on the stochastic perturbations and control sequences. Section 4 presents the main convergence results: Lemma 4.1 establishes mean-square boundedness of the iterates, Theorem 4.2 proves almost sure convergence of the entire sequence. Finally, Section 5 summarizes the contributions and discusses some avenues for future research, including adaptive variants and empirical benchmarking.
2 Related Work and Contributions
The Halpern iteration originated as a fixed-point method for nonexpansive mappings and, under standard control conditions on the weight sequence, guarantees strong convergence toward the anchor-selected fixed point, namely the metric projection of the anchor onto the fixed-point set, see e.g. [23, 42, 43].
Recent works have clarified sharp non-asymptotic residual guarantees for Halpern-type schemes. Early non-asymptotic $O(1/k)$ residual bounds for viscosity/Halpern-type iterations can be found e.g. in [39]. More recently, worst-case analyses have identified (near-)minimax optimal parameter choices and tight bounds for broad classes of Mann iterations; see [15, 9]. Related optimal-complexity and matching lower-bound results for fixed-point problems are developed in [32].
In the context of monotone inclusions and variational inequalities, [16] shows how variants of the Halpern iteration can produce near-optimal and parameter-free methods for standard classes of problems with Lipschitz monotone operators.
These initial results are formulated in an essentially deterministic framework; as for the stochastic case, the available works tend to impose variance control, obtained often via increasing mini-batches or variance reduction, to prevent error accumulation.
A starting point is [12], which studies stochastic monotone inclusions and combines variants of Halpern with a recursive PAGE-type estimator to obtain a mean-square error (variance) bound for the estimator that decreases as $O(\varepsilon^{2}/k)$, via data-driven choices of the batch sizes. Anyway he work assumes a bounded variace and a multi-point oracle assumption, which allows to query the stochastic oracle at multiple points under common randomness (i.e., using the same sample/seed). property in expectation. This line therefore puts at its core the idea that, with constant-variance noise, additional mechanisms are needed to make the effective error decay [12].
A related perspective, closer to fixed-point theory, is developed in [11], which analyzes a stochastic Halpern iteration in normed spaces and discusses applications in reinforcement learning. Here, the stochastic oracle is modeled as a random approximation $\widetilde{T}(x,\xi)$ of a nonexpansive operator $T$ and the almost sure convergence to a fixed point under the (rather strong) summability condition $\sum_{n}\mathbb{E}\|U_{n}\|<+\infty,$ where $\{U_{n}\}$ is the error sequence (Theorem 3.10).
Related “anchoring” ideas are also leveraged to obtain near-optimal sample complexity in reinforcement learning: [27] proposes an anchored (Halpern-type) approach for computing $\varepsilon$-optimal policies in MDPs, with near-optimal sample and time complexity guarantees. In this direction, [33] proposes a general framework of stochastic Halpern scheme for nonexpansive operators, where stochasticity enters as additive perturbations $\xi_{n},\delta_{n}.$ Here, the primary objective is to establish the asymptotic regularity of the sequence of iterates, that is, the convergence $\|x_{n+1}-x_{n}\|\to 0$.
This type of result is obtained under standard control conditions on $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ (e.g., including bounded-variation requirements), together with suitable summability assumptions on the perturbation increments, such as $\sum_{n}\mathbb{E}\|\delta_{n+1}-\delta_{n}\|<+\infty$ and $\sum_{n}\mathbb{E}\|\xi_{n+1}-\xi_{n}\|<+\infty$, under a mild boundedness-in-expectation assumption on the iterates.
Alongside the works specifically focused on Halpern-type schemes, general results on stochastic fixed-point iterations are also relevant. In particular, [10] analyzes Stochastic Krasnosel’skiĭ–Mann (SKM) iterations for nonexpansive operators, where the noise sequence is modeled as additive stochastic perturbations with zero mean and finite second moment, that is, $\mathbb{E}[U_{n}\mid\mathcal{F}_{n-1}]=0$ and $\mathbb{E}\|U_{n}\|^{2}<\infty$. Under standard stepsize assumptions, the authors establish almost sure convergence of the SKM iterates to a fixed point of $T$, without requiring the variance to be uniformly bounded.
In our work we place ourselves in this latter perspective. Indeed, we study HalpernSGD as an “anchored” scheme for $L$-smooth convex minimization, modeling gradients stochastically perturbed by an unbiased noise sequence. Almost sure convergence in our setting is established in a different way than in the works reviewed above. Specifically, we study almost sure convergence to a minimizer, that is, to a zero of the gradient of the objective function $f$, by introducing a second control sequence $\{\epsilon_{n}\}$ that plays the role of a learning rate. This additional degree of control allows us to impose only a mild condition on the noise, namely the summability condition
$\sum_{n}\epsilon_{n}^{2}\,\|U_{n}\|^{2}<+\infty\qquad\text{a.s.},$
which, under the bounded-variance assumption and $\sum_{n=0}^{\infty}\epsilon_{n}^{2}<\infty$, holds automatically almost surely and ensures that the cumulative effect of the stochastic perturbations remains tractable. As a consequence, we are able to handle almost sure convergence of the entire sequence $\{x_{n}\}$ even in the presence of noise with merely bounded variance. This, in turn, allows the Halpern framework to be applied in settings where variance reduction or vanishing-noise assumptions are not available.
By exploiting a quadratic-mean recurrence and lemma proved by Xu in [43, 44], we first establish mean-square convergence, i.e., $\mathbb{E}\|x_{n}-x^{*}\|^{2}\to 0$ , and we then strengthen the result to almost sure convergence of the entire sequence to the anchor-selected minimizer $x^{*}=P_{S}(u)$.
Convergence-rate analyses for Halpern-type stabilized stochastic gradient schemes, and connections to variance growth conditions beyond the classical uniformly bounded-variance model, have recently been developed in [31, 1]; see in particular Section 3 of [1].
In summary, existing stochastic analyses of Halpern-type schemes typically rely either on (i) variance reduction / increasing-batch mechanisms to enforce a decreasing effective noise level [12], or on (ii) summability/vanishing conditions on the perturbations to prevent error accumulation [11, 33]. General stochastic fixed-point results for nonexpansive maps (not necessarily Halpern-type) are also relevant; see [10]. Our contribution aims at an almost sure guarantee of convergence under bounded-variance assumptions and is based on an appropriate choice and coupling of two control sequences that make the cumulative effect of the noise tractable.
3 Preliminaries
Throughout, $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ denote the Euclidean inner product and norm on $\mathbb{R}^{d}$. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and differentiable with $L$-Lipschitz continuous gradient, i.e.,
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\qquad\forall x,y\in\mathbb{R}^{d}.$ (5)
We denote the set of minimizers by
$S:=\arg\min_{x\in\mathbb{R}^{d}}f(x),$ (6)
and assume $S\neq\emptyset$. Given an anchor $u\in\mathbb{R}^{d}$, we write $P_{S}(u)$ for the (unique) metric projection of $u$ onto $S$ (recall that $S$ is closed and convex for convex $f$).
We model stochastic gradients via a martingale-difference noise sequence $\{U_{n}\}_{n\geq 0}$, whose assumptions will be stated explicitly later on. Denoting the step sizes by $\{\epsilon_{n}\}_{n\geq 0}\subset(0,\infty)$ and the Halpern weights by $\{\alpha_{n}\}_{n\geq 0}\subset(0,1)$, we define the stochastic forward (gradient) operator as
$\widetilde{T}_{n}(x):=x-\epsilon_{n}\bigl(\nabla f(x)+U_{n}\bigr),$ (7)
and the HalpernSGD iteration reads
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\,\widetilde{T}_{n}(x_{n})=\alpha_{n}u+(1-\alpha_{n})\Bigl(x_{n}-\epsilon_{n}\bigl(\nabla f(x_{n})+U_{n}\bigr)\Bigr).$ (8)
When $U_{n}\equiv 0$, (8) reduces to the deterministic Halpern iterative scheme. In this deterministic regime, strong convergence is established in [20] under the following coupled conditions
  • $\lim_{n\to\infty}\alpha_{n}=0,\quad\sum_{n=0}^{\infty}\alpha_{n}=+\infty,\quad\sum_{n=0}^{\infty}\epsilon_{n}^{\,2}<+\infty,$
  • $\lim_{n\to\infty}\dfrac{\epsilon_{n}}{\alpha_{n}}=0,\quad\sum_{n=0}^{\infty}\epsilon_{n}\alpha_{n}<+\infty.$
In the stochastic context, a decreasing stepsize sequence $\{\epsilon_{n}\}$ is required to balance progress and noise reduction when moving from deterministic to stochastic gradients. Classical stochastic approximation theory [37] prescribes the following conditions:
$\sum_{n=1}^{\infty}\epsilon_{n}=\infty,\qquad\sum_{n=1}^{\infty}\epsilon_{n}^{2}<\infty.$ (9)
As it will be specified in the next section, in our analysis, $\{\epsilon_{n}\}$ will be additionally coupled with $\{\alpha_{n}\}$ through Halpern-type control conditions.
We recall a few standard facts that will be used in the next section. Convexity of $f$ implies monotonicity of the gradient:
$\langle\nabla f(x)-\nabla f(y),\,x-y\rangle\geq 0\qquad\forall x,y\in\mathbb{R}^{d}.$ (10)
Moreover, the $L$-smoothness assumption (5) together with convexity yields the Baillon–Haddad cocoercivity inequality
$\langle\nabla f(x)-\nabla f(y),\,x-y\rangle\geq\frac{1}{L}\,\|\nabla f(x)-\nabla f(y)\|^{2}\qquad\forall x,y\in\mathbb{R}^{d}.$ (11)
In particular, if $x^{*}\in S$, then $\nabla f(x^{*})=0$ and (11) gives
$\langle\nabla f(x),\,x-x^{*}\rangle\geq\frac{1}{L}\,\|\nabla f(x)\|^{2}\qquad\forall x\in\mathbb{R}^{d}.$ (12)
We also recall the standard smooth-convex gradient lower bound:
$\frac{1}{2L}\,\|\nabla f(x)\|^{2}\leq f(x)-f^{*}\qquad\forall x\in\mathbb{R}^{d},$ (13)
where $f^{*}:=\min_{\mathbb{R}^{d}}f$. Finally, since $f$ is convex and differentiable, minimizers are characterized by the first-order condition
$x^{*}\in S\quad\Longleftrightarrow\quad\nabla f(x^{*})=0.$ (14)
4 Convergence Analysis of HalpernSGD
We now investigate the convergence properties of the HalpernSGD iteration (8). Throughout this section, we consider the standard stochastic setting where $\{\mathcal{F}_{n}\}$ denotes the natural filtration generated by the iterates and the random perturbations, i.e.
$\mathcal{F}_{n}:=\sigma(x_{0},U_{0},\dots,U_{n-1}).$
The stepsizes $\{\epsilon_{n}\}$ and Halpern weights $\{\alpha_{n}\}$ are real sequences satisfying
$\epsilon_{n}>0,\quad\epsilon_{n}\to 0,\quad\sum_{n=0}^{\infty}\epsilon_{n}=+\infty,\quad\sum_{n=0}^{\infty}\epsilon_{n}^{2}<+\infty,$
and
$\alpha_{n}\in(0,1),\quad\alpha_{n}\to 0,\quad\sum_{n=0}^{\infty}\alpha_{n}=+\infty,\quad\sum_{n=0}^{\infty}\alpha_{n}^{2}<+\infty$
with
$\lim_{n\rightarrow+\infty}\frac{\epsilon_{n}}{\alpha_{n}}=+\infty\quad\text{and }\sum_{n=0}^{\infty}\alpha_{n}\epsilon_{n}<\infty.$
Remark 1 (A concrete admissible choice of $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$).
We stress that the above coupled conditions on the Halpern weights and step sizes are not particularly restrictive. For instance, the one-parameter choice
$\alpha_{n}:=\frac{1}{n+1},\qquad\epsilon_{n}:=\frac{1}{(n+1)^{p}},\qquad p\in\Bigl(\tfrac{1}{2},1\Bigr),$
satisfies all the stated assumptions. In addition, this example also satisfies $\lim_{n\to\infty}\frac{\alpha_{n}}{\epsilon_{n}^{2}}=+\infty,$ which is a condition required at a later stage of our convergence analysis.
As for the stochastic perturbations $U_{n}$, we assume the standard martingale difference condition (see, e.g., [10])
$\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0\quad\text{a.s.}$
In place of the classical uniformly bounded-variance assumption, we impose the summability condition
$\sum_{n=0}^{\infty}\epsilon_{n}^{2}\,\mathbb{E}\!\left[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\right]<+\infty\quad\text{a.s.}$
This requirement is strictly weaker than the standard bounded conditional variance hypothesis. For a deeper discussions, we refer to [1].
Lemma 4.1.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a convex and $L$-smooth function, and let $\{x_{n}\}$ be the sequence generated by the HalpernSGD iteration (8). Let $x^{*}\in S$ (equivalently, $\nabla f(x^{*})=0$) and assume that the stochastic perturbations and the parameters $\{\alpha_{n}\},\{\epsilon_{n}\}$ satisfy the conditions stated above, and in addition that $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$. Then, the sequence $\{\mathbb{E}\|x_{n}-x^{*}\|^{2}\}_{n\in\mathbb{N}}$ is bounded.
Proof.
Let $x^{*}\in S$, i.e. $\nabla f(x^{*})=0$. By definition of the HalpernSGD iteration, we have
$x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\big(x_{n}-\epsilon_{n}(\nabla f(x_{n})+U_{n})\big),$
which yields
$\|x_{n+1}-x^{*}\|^{2}=\big\|\alpha_{n}(u-x^{*})+(1-\alpha_{n})\big(x_{n}-x^{*}-\epsilon_{n}(\nabla f(x_{n})+U_{n})\big)\big\|^{2}.$
Let us denote $g_{n}(x_{n})=\nabla f(x_{n})+U_{n}$. Using the convexity of the squared norm, we obtain
$\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\|x_{n}-x^{*}-\epsilon_{n}g_{n}(x_{n})\|^{2},$
and expanding the second term gives
$\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\!\left[\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\|g_{n}(x_{n})\|^{2}-2\epsilon_{n}\langle g_{n}(x_{n}),x_{n}-x^{*}\rangle\right].$
Taking the conditional expectation with respect to $\mathcal{F}_{n}$ yields
$\mathbb{E}\!\left[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}\right]\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})\epsilon_{n}\langle\mathbb{E}[g_{n}(x_{n})\mid\mathcal{F}_{n}],x_{n}-x^{*}\rangle$
$(1-\alpha_{n})\epsilon_{n}^{2}\mathbb{E}\!\left[\|g_{n}(x_{n})\|^{2}\mid\mathcal{F}_{n}\right].$
We analyze each term separately.
  1. 1.
    Since $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$, we have $\mathbb{E}[g_{n}(x_{n})\mid\mathcal{F}_{n}]=\nabla f(x_{n})$. Using convexity of $f$, it follows that
    $\langle\nabla f(x_{n})-\nabla f(x^{*}),x_{n}-x^{*}\rangle\geq 0,$
    so that the cross term is nonpositive:
    $2(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle\leq 0.$
  2. 2.
    Taking total expectations on both sides yields the corresponding inequality for $\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$.
  3. 3.
    Using $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$ again, we have that
    $\mathbb{E}\left[\langle U_{n},\nabla f(x_{n})\rangle\right]=\mathbb{E}\Big[\mathbb{E}\big[\langle U_{n},\nabla f(x_{n})\rangle\mid\mathcal{F}_{n}\big]\Big]=\mathbb{E}\Big[\langle\mathbb{E}\big[U_{n}\mid\mathcal{F}_{n}\big],\nabla f(x_{n})\rangle\Big]=0\text{ and}$
    $\mathbb{E}\|g_{n}(x_{n})\|^{2}=\mathbb{E}\|\nabla f(x_{n})+U_{n}-\nabla f(x^{*})\|^{2}=\mathbb{E}\|\nabla f(x_{n})-\nabla f(x^{*})\|^{2}+\mathbb{E}\|U_{n}\|^{2}.$
    Since $f$ is $L$-smooth, $\|\nabla f(x_{n})-\nabla f(x^{*})\|\leq L\|x_{n}-x^{*}\|$, hence
    $\mathbb{E}\|g_{n}(x_{n})\|^{2}\leq L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\mathbb{E}\|U_{n}\|^{2}.$
Combining these estimates and taking expectations, we obtain
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\bigl(\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\mathbb{E}\|U_{n}\|^{2}\bigr)$
$\leq\alpha_{n}\|u-x^{*}\|^{2}+(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}L^{2}\mathbb{E}\|x_{n}-x^{*}\|^{2}+\epsilon_{n}^{2}\mathbb{E}\|U_{n}\|^{2}.$
Let $a_{n}:=\mathbb{E}\|x_{n}-x^{*}\|^{2}$ and $\mu_{n}:=\mathbb{E}\|U_{n}\|^{2}$. Then
$a_{n+1}\leq\alpha_{n}\|u-x^{*}\|^{2}+\big((1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\big)a_{n}+\epsilon_{n}^{2}\mu_{n}.$
$(\alpha_{n}-\epsilon_{n}^{2}L^{2})\|u-x^{*}\|^{2}+\big[1-(\alpha_{n}-\epsilon_{n}^{2}L^{2})\big]a_{n}+\epsilon_{n}^{2}\mu_{n}+\epsilon_{n}^{2}L^{2}\|u-x^{*}\|^{2}.$
Setting $\gamma_{n}:=\alpha_{n}-\epsilon_{n}^{2}L^{2}$, we can rewrite
$a_{n+1}\leq\gamma_{n}\|u-x^{*}\|^{2}+(1-\gamma_{n})a_{n}+\epsilon_{n}^{2}(\mu_{n}+L^{2}\|u-x^{*}\|^{2}).$
We remark that $\gamma_{n}\in[0,1)$. Indeed, $\gamma_{n}=\alpha_{n}-\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, and by the coupling condition $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}$, we have $\gamma_{n}\geq 0$.
Then, iterating the inequality gives
$a_{n+1}\leq\max\{\|u-x^{*}\|^{2},a_{0}\}+\sum_{k=0}^{n}\epsilon_{k}^{2}\big[\mu_{k}+L^{2}\|u-x^{*}\|^{2}\big].$
Now, by assumptions, $\sum_{k}\epsilon_{k}^{2}<\infty$ and the stochastic perturbations satisfy $\mathbb{E}[\|U_{k}\|^{2}\mid\mathcal{F}_{k}]\leq\sigma^{2}$ a.s. $\forall k\in\mathbb{N}$, then $\mu_{k}\leq\sigma^{2}$ and
$\sum_{k=0}^{\infty}\epsilon_{k}^{2}(\mu_{k}+L^{2}\|u-x^{*}\|^{2})\leq(\sigma^{2}+L^{2}\|u-x^{*}\|^{2})\sum_{k=0}^{\infty}\epsilon_{k}^{2}<\infty.$
Hence, the sequence $\{a_{n}\}$ is bounded, that is, $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty.$
Remark 2.
From a practical standpoint, in view of the experimental analysis, the nonnegativity condition $\alpha_{n}-\epsilon_{n}^{2}L^{2}\geq 0$ (equivalently, $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$) can be ensured through a simple sufficient asymptotic requirement.
Indeed, by assuming
$\lim_{n\to+\infty}\frac{\alpha_{n}}{\epsilon_{n}^{2}}=+\infty,$ (15)
it follows that there exists $\tilde{n}\in\mathbb{N}$ such that
$\forall n>\tilde{n},\qquad\frac{\alpha_{n}}{\epsilon_{n}^{2}}>L^{2},$
which directly ensures $\alpha_{n}-\epsilon_{n}^{2}L^{2}>0$ for all sufficiently large $n$. This condition is easier to verify in numerical implementations, where both sequences $\{\alpha_{n}\}$ and $\{\epsilon_{n}\}$ are explicitly specified.
Theorem 4.2 (Mean-square convergence).
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth, $S:=\arg\min f\neq\emptyset$, $x^{*}:=P_{S}(u)$, and let $\{x_{n}\}$ be generated by (8). Assume the stochastic setting and the step-size conditions stated above, together with
$(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1\qquad\forall n.$
Then
$\mathbb{E}\|x_{n}-x^{*}\|^{2}\longrightarrow 0.$
In particular, $x_{n}\to x^{*}$ in probability, and there exists a subsequence $(x_{n_{k}})$ such that $x_{n_{k}}\to x^{*}$ almost surely. Moreover, $x^{*}$ is the unique point in $S$ satisfying the variational inequality
$\langle u-x^{*},\,p-x^{*}\rangle\leq 0\qquad\text{for all }p\in S.$
Proof.
We show that $\mathbb{E}\|x_{n}-x^{*}\|^{2}\to 0$ under the stated assumptions. By definition of HalpernSGD iteration, we have
$\|x_{n+1}-x^{*}\|^{2}=\bigl\|\alpha_{n}(u-x^{*})+(1-\alpha_{n})\bigl(x_{n}-x^{*}-\epsilon_{n}g_{n}\bigr)\bigr\|^{2}\qquad g_{n}:=\nabla f(x_{n})+U_{n}.$
Expanding the square we obtain:
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,x_{n}-x^{*}-\epsilon_{n}g_{n}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}-\epsilon_{n}g_{n}\|^{2}$
$\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle g_{n},u-x^{*}\rangle$
$\qquad+(1-\alpha_{n})^{2}\bigl(\|x_{n}-x^{*}\|^{2}-2\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+\epsilon_{n}^{2}\|g_{n}\|^{2}\bigr).$
Collecting terms, we can rewrite this as
$\|x_{n+1}-x^{*}\|^{2}=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle g_{n},u-x^{*}\rangle$
$\qquad+(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}-2(1-\alpha_{n})^{2}\epsilon_{n}\langle g_{n},x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\|g_{n}\|^{2}.$
Now, taking the conditional expectation with respect to $\mathcal{F}_{n}$ and using $\mathbb{E}[g_{n}\mid\mathcal{F}_{n}]=\nabla f(x_{n})$ (since $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$), we obtain
$\mathbb{E}\!\left[\|x_{n+1}-x^{*}\|^{2}\mid\mathcal{F}_{n}\right]$ $=\alpha_{n}^{2}\|u-x^{*}\|^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},x_{n}-x^{*}\rangle$
$\quad-2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\langle\nabla f(x_{n}),u-x^{*}\rangle+(1-\alpha_{n})^{2}\|x_{n}-x^{*}\|^{2}$
$\quad-2(1-\alpha_{n})^{2}\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\,\mathbb{E}\!\left[\|g_{n}\|^{2}\mid\mathcal{F}_{n}\right].$
First, since $f$ is convex and $L$-smooth, by the co-coercivity inequality we have
$2(1-\alpha_{n})^{2}\epsilon_{n}\langle\nabla f(x_{n}),x_{n}-x^{*}\rangle\;\leq\;-\frac{2(1-\alpha_{n})^{2}\epsilon_{n}}{L}\,\|\nabla f(x_{n})\|^{2}.$
Next, from Lemma 4.1, there exists $C>0$ such that $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\leq C$. Moreover, by $L$-Lipschitz continuity of $\nabla f$ and $\nabla f(x^{*})=0$,
$\|\nabla f(x_{n})\|\leq L\|x_{n}-x^{*}\|.$
Hence, by Cauchy–Schwarz,
$\mathbb{E}\big|\langle\nabla f(x_{n}),u-x^{*}\rangle\big|\leq\|u-x^{*}\|\,\mathbb{E}\|\nabla f(x_{n})\|\leq L\|u-x^{*}\|\,\mathbb{E}\|x_{n}-x^{*}\|\leq L\sqrt{C}\,\|u-x^{*}\|.$
Therefore,
$2\alpha_{n}(1-\alpha_{n})\epsilon_{n}\,\mathbb{E}\langle\nabla f(x_{n}),u-x^{*}\rangle\leq 2\alpha_{n}(1-\alpha_{n})\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
Finally, by the martingale-difference assumption and bounded conditional variance,
$\mathbb{E}\!\left[\|g_{n}\|^{2}\mid\mathcal{F}_{n}\right]=\|\nabla f(x_{n})\|^{2}+\mathbb{E}\!\left[\|U_{n}\|^{2}\mid\mathcal{F}_{n}\right]\leq\|\nabla f(x_{n})\|^{2}+\sigma^{2}.$
Then, taking the total expectation and using $(1-\alpha_{n})\leq 1$, we obtain
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$ $\leq\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}-(1-\alpha_{n})^{2}\!\Bigl(\tfrac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\,\mathbb{E}\|\nabla f(x_{n})\|^{2}$
$\quad+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$ (16)
Set $a_{n}:=\mathbb{E}\|x_{n}-x^{*}\|^{2}$ and define
$\kappa_{n}:=\frac{(1-\alpha_{n})^{2}}{\alpha_{n}}\Bigl(\frac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\geq 0,\qquad b_{n}:=2(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle-\kappa_{n}\,\mathbb{E}\|\nabla f(x_{n})\|^{2},$
$c_{n}:=\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$
Then (16) can be rewritten as
$a_{n+1}\leq(1-\alpha_{n})a_{n}+\alpha_{n}b_{n}+c_{n}.$ (17)
By the step-size assumptions, $\sum_{n}\alpha_{n}^{2}<\infty$, $\sum_{n}\epsilon_{n}^{2}<\infty$ and $\sum_{n}\alpha_{n}\epsilon_{n}<\infty$, hence $\sum_{n}c_{n}<\infty$.
Step 1: $\limsup_{n\to\infty}b_{n}\leq 0$. First note that, by Lemma 4.1, there exists $C_{0}>0$ such that $\sup_{n}a_{n}\leq C_{0}$. Hence, by Cauchy–Schwarz,
$\big|2(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle\big|\leq 2\|u-x^{*}\|\,\sqrt{a_{n}}\leq 2\|u-x^{*}\|\sqrt{C_{0}}=:M.$ (18)
Moreover, by the coupling $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ we have $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, hence $\epsilon_{n}\leq 1/L$ for all $n$. Therefore
$\Bigl(\frac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)=\epsilon_{n}\Bigl(\frac{2}{L}-\epsilon_{n}\Bigr)\geq\frac{\epsilon_{n}}{L}.$
Therefore,
$\kappa_{n}\;\geq\;\frac{(1-\alpha_{n})^{2}}{\alpha_{n}}\cdot\frac{\epsilon_{n}}{L}\;\xrightarrow[n\to\infty]{}\;+\infty\qquad\text{since }\frac{\epsilon_{n}}{\alpha_{n}}\to+\infty.$ (19)
We claim that $\limsup_{n\to\infty}b_{n}\leq 0$. Suppose by contradiction that $\limsup_{n\to\infty}b_{n}=\delta>0$. Then there exists a subsequence $(n_{k})$ such that $b_{n_{k}}\to\delta$ and hence $b_{n_{k}}\geq\delta/2$ for $k$ large enough. Using (18),
$\kappa_{n_{k}}\,\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}=2(1-\alpha_{n_{k}})\langle u-x^{*},\mathbb{E}[x_{n_{k}}-x^{*}]\rangle-b_{n_{k}}\leq M-\frac{\delta}{2},$
so, by (19),
$\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\longrightarrow 0.$ (20)
In particular, by Markov’s inequality, $\nabla f(x_{n_{k}})\to 0$ in probability.
Since $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty$, the family $\{x_{n}\}$ is tight.
Thus, passing to a further subsequence (not relabeled), we may assume that the laws of $x_{n_{k}}$ converge weakly to some probability measure $\mu$ on $\mathbb{R}^{d}$. Equivalently, there exists an $\mathbb{R}^{d}$-valued random vector $P$ (defined on some probability space) with law $\mu$ such that
$x_{n_{k}}\Rightarrow P\qquad\text{in distribution.}$
By the continuous mapping theorem, $\nabla f(x_{n_{k}})\Rightarrow\nabla f(P)$. On the other hand, $\nabla f(x_{n_{k}})\to 0$ in probability implies $\nabla f(x_{n_{k}})\Rightarrow 0$, hence $\nabla f(P)=0$ almost surely, i.e. $P\in S$ almost surely.
Now define $X_{n}:=\langle x_{n}-x^{*},\,u-x^{*}\rangle$. By Cauchy–Schwarz, $\sup_{n}\mathbb{E}|X_{n}|^{2}\leq\|u-x^{*}\|^{2}\sup_{n}a_{n}<\infty$, hence $\{X_{n}\}$ is uniformly integrable.
Since $x_{n_{k}}\Rightarrow P$, we have $X_{n_{k}}\Rightarrow X:=\langle P-x^{*},u-x^{*}\rangle$. Since $\{X_{n_{k}}\}$ is uniformly integrable and $X_{n_{k}}\Rightarrow X$, we have
$\mathbb{E}X_{n_{k}}\longrightarrow\mathbb{E}X.$
Since $P(\omega)\in S$ a.s. and $x^{*}=P_{S}(u)$, the projection characterization gives $\langle u-x^{*},P-x^{*}\rangle\leq 0$ a.s., hence $\mathbb{E}\langle P-x^{*},u-x^{*}\rangle\leq 0$. Therefore,
$\limsup_{k\to\infty}2(1-\alpha_{n_{k}})\langle u-x^{*},\mathbb{E}[x_{n_{k}}-x^{*}]\rangle\leq 0.$
Together with (20) and $\kappa_{n_{k}}\geq 0$, this implies
$\limsup_{k\to\infty}b_{n_{k}}\leq 0,$
contradicting $b_{n_{k}}\to\delta>0$. Hence $\limsup_{n\to\infty}b_{n}\leq 0$.
Step 2: apply Xu’s lemma. We have shown that (17) holds with $\sum_{n}c_{n}<\infty$ and $\limsup b_{n}\leq 0$. Since also $\sum_{n}\alpha_{n}=\infty$ and $\alpha_{n}\in(0,1)$, Xu’s lemma (see [43, Lemma 2.1]) yields $a_{n}\to 0$, i.e.
$\mathbb{E}\|x_{n}-x^{*}\|^{2}\longrightarrow 0.$
Convergence in probability follows from Chebyshev’s inequality. Finally, choose integers $n_{k}\uparrow\infty$ such that $\mathbb{E}\|x_{n_{k}}-x^{*}\|^{2}\leq 2^{-k}$. Then, for any $\nu>0$, by Markov,
$\sum_{k=1}^{\infty}\mathbb{P}\big(\|x_{n_{k}}-x^{*}\|>\nu\big)\leq\frac{1}{\nu^{2}}\sum_{k=1}^{\infty}\mathbb{E}\|x_{n_{k}}-x^{*}\|^{2}\leq\frac{1}{\nu^{2}}\sum_{k=1}^{\infty}2^{-k}<\infty.$
By Borel–Cantelli, $\|x_{n_{k}}-x^{*}\|\to 0$ almost surely. ∎
The following lemma is a standard consequence of Baillon–Haddad cocoercivity; see, e.g., [3, 4].
Lemma 4.3 (Nonexpansiveness of a gradient step).
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth. Then for every $\epsilon\in[0,2/L]$, the mapping $T_{\epsilon}(x):=x-\epsilon\nabla f(x)$ is nonexpansive:
$\|T_{\epsilon}(x)-T_{\epsilon}(y)\|\leq\|x-y\|\qquad\forall x,y\in\mathbb{R}^{d}.$
Proof.
Using cocoercivity (11), for all $x,y$ we have
$\|T_{\epsilon}(x)-T_{\epsilon}(y)\|^{2}=\|x-y-\epsilon(\nabla f(x)-\nabla f(y))\|^{2}$
$\|x-y\|^{2}-2\epsilon\langle\nabla f(x)-\nabla f(y),x-y\rangle+\epsilon^{2}\|\nabla f(x)-\nabla f(y)\|^{2}$
$\leq\|x-y\|^{2}-\Bigl(\frac{2\epsilon}{L}-\epsilon^{2}\Bigr)\|\nabla f(x)-\nabla f(y)\|^{2}\leq\|x-y\|^{2},$
whenever $\epsilon\leq 2/L$. Taking square roots gives the claim. ∎
Theorem 4.4 (Almost sure convergence).
Assume the hypotheses of Theorem 4.2. Then
$x_{n}\longrightarrow x^{*}\qquad\text{almost surely.}$
Proof.
Step 0: define the noise-free reference sequence. Let $\{\bar{x}_{n}\}$ be the noise-free Halpern gradient sequence defined by
$\bar{x}_{n+1}:=\alpha_{n}u+(1-\alpha_{n})\bigl(\bar{x}_{n}-\epsilon_{n}\nabla f(\bar{x}_{n})\bigr),\qquad\bar{x}_{0}:=x_{0}.$
For each fixed initial point $\bar{x}_{0}\in\mathbb{R}^{d}$, the deterministic recursion above converges to $x^{*}=P_{S}(u)$ (this is the deterministic special case of Theorem 4.2, obtained by setting $U_{n}\equiv 0$; since the recursion is deterministic, the conclusion is pointwise). Hence, for the (possibly random) initial $x_{0}$, we still have
$\bar{x}_{n}(\omega)\longrightarrow x^{*}\qquad\text{for all }\omega\in\Omega.$ (21)
Step 1: the stochastic iterates track the noise-free iterates almost surely. Define the error $e_{n}:=x_{n}-\bar{x}_{n}$. Using the updates for $x_{n+1}$ and $\bar{x}_{n+1}$, we get
$e_{n+1}=(1-\alpha_{n})\Bigl((x_{n}-\epsilon_{n}\nabla f(x_{n}))-(\bar{x}_{n}-\epsilon_{n}\nabla f(\bar{x}_{n}))-\epsilon_{n}U_{n}\Bigr)=(1-\alpha_{n})\bigl(T_{\epsilon_{n}}(x_{n})-T_{\epsilon_{n}}(\bar{x}_{n})-\epsilon_{n}U_{n}\bigr),$
where $T_{\epsilon}(x)=x-\epsilon\nabla f(x)$.
Taking squared norms and conditional expectation with respect to $\mathcal{F}_{n}$, and using that $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$, we obtain
$\mathbb{E}\bigl[\|e_{n+1}\|^{2}\mid\mathcal{F}_{n}\bigr]$ $=(1-\alpha_{n})^{2}\Bigl(\|T_{\epsilon_{n}}(x_{n})-T_{\epsilon_{n}}(\bar{x}_{n})\|^{2}+\epsilon_{n}^{2}\,\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\Bigr).$
Note that $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ implies $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, hence $\epsilon_{n}\leq 1/L<2/L$, so Lemma 4.3 applies to $T_{\epsilon_{n}}$. By Lemma 4.3 (nonexpansiveness of $T_{\epsilon_{n}}$) and by $\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\leq\sigma^{2}$, we get
$\mathbb{E}\bigl[\|e_{n+1}\|^{2}\mid\mathcal{F}_{n}\bigr]\leq(1-\alpha_{n})^{2}\|e_{n}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}\leq(1-\alpha_{n})\|e_{n}\|^{2}+\epsilon_{n}^{2}\sigma^{2},$
where we used $(1-\alpha_{n})^{2}\leq(1-\alpha_{n})$ and $(1-\alpha_{n})^{2}\leq 1$.
Set $V_{n}:=\|e_{n}\|^{2}$. Then
$\mathbb{E}[V_{n+1}\mid\mathcal{F}_{n}]\leq V_{n}-\alpha_{n}V_{n}+\epsilon_{n}^{2}\sigma^{2}.$
Since $\sum_{n}\epsilon_{n}^{2}<\infty$, the Robbins–Siegmund almost-supermartingale theorem (see, e.g., [18]) implies that $V_{n}$ converges almost surely and
$\sum_{n=0}^{\infty}\alpha_{n}V_{n}<\infty\qquad\text{a.s.}$
Because $\sum_{n}\alpha_{n}=\infty$ and $V_{n}\geq 0$, the latter forces $\liminf_{n\to\infty}V_{n}=0$. Since $V_{n}$ also converges a.s., we conclude that
$\|e_{n}\|^{2}=V_{n}\longrightarrow 0\qquad\text{a.s.}$ (22)
Step 2: conclude. By (21) and (22),
$\|x_{n}-x^{*}\|\leq\|x_{n}-\bar{x}_{n}\|+\|\bar{x}_{n}-x^{*}\|\longrightarrow 0\qquad\text{a.s.}$
Hence $x_{n}\to x^{*}$ almost surely. ∎
5 Conclusion and future perspectives
In this work we addressed the convergence theory for the HalpernSGD optimizer. Building on tools from smooth convex analysis and stochastic approximation, we proved that its iterates converge almost surely to a specific solution selected by the anchor, namely $x^{*}=P_{S}(u)$.
As future work, we plan to broaden the experimental analysis carried out in [19, 14], where HalpernSGD was compared mainly with SGD under constant learning rates, in order to assess whether HalpernSGD can be among the most efficient optimizers within the gradient-based family while maintaining (or improving) accuracy. In particular, we aim to benchmark HalpernSGD against SGD [37], SGD with momentum [40], NAG [30], and HB [34] and other accelerated gradient schemes. Motivated by the empirical benefits observed for the anchoring mechanism in [19, 14], we also plan to investigate anchored variants of adaptive methods. A natural direction is to combine an anchor with an Adam-type update through a convex blending (an “Anchored-Adam” scheme), analogously to HalpernSGD as an anchored variant of SGD. Preliminary experiments suggest that Halpern-style anchoring may act as a general acceleration and stabilization mechanism across different optimizer families. Establishing convergence guarantees for such anchored adaptive schemes remains an open theoretical challenge, especially in view of the known subtle convergence behavior of Adam optimizer, and we consider this a key direction for future research.
Funding
This publication was partially funded by the PhD program in Mathematics and Computer Science at University of Calabria, Cycle XXXVIII with the support of a scholarship financed by DM 351/2022 (CUP H23C22000440007), based on the NRRP funded by the European Union.